6.Factory Overhead Cost Variances

The following data relate to factory overhead cost for the production of 4,000 computers:

Actual: Variable factory overhead $74,700
Fixed factory overhead 28,500
Standard: 4,000 hrs. at $24 96,000

If productive capacity of 100% was 6,000 hours and the factory overhead cost budgeted at the level of 4,000 standard hours was $105,500, determine the variable factory overhead Controllable Variance, fixed factory overhead volume variance, and total factory overhead cost variance. The fixed factory overhead rate was $4.75 per hour. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variance Amount Favorable/Unfavorable
Controllable variance $ Favorable
Volume variance $ Unfavorable
Total factory overhead cost variance: $ Unfavorable
Answer #1

a)

Calculation of Variable Factory Overhead controllable Variance:

Variable Factory OH controllable Variance = Actual variable Factory OH - Variable OH for Standard Hours allowed.

Actual variable Factory OH = Total Budgeted Factory OH-Budgeted Fixed Factory

= 74,700 - 105,500-(6,000*4.75)

= 74,700 - 105,500-28,500

= 74,700 - 77,000

= 2,300 F

b)

Calculation of Fixed Factory Volume Variance:

Fixed Factory Volume Variance = Budgeted fixed OH - Applied fixed OH

Budgeted fixed OH = Standard labor hours for Actual production*standard Rate

= 6,000*4.75 - 4,000*4.75

= 28,500-19,000

= 9,500 U

c)

Calculation of Total Factory Overhead Cost Variance:

Total Factory OH cost variance = Controllable Variance+Volume Variance

-2,300+9,500

= 7,200 U